OpenStudy (anonymous):

Sharon sails for 6 miles at a speed of 3 mph. She then jumps in the water and swims at a speed of 2 mph for 2 miles. For how long does Sharon travel?

OpenStudy (wolfe8):

Use the relationship v=d/t or rearranged t=d/v